OpenStudy (anonymous):

Given an 8-bit binary number/counter, why is it that 11111111 converts to -1 in decimal format?

OpenStudy (anonymous):

Good question =) Well notice that there are different methods for that. It will be -1 in two's complement system. Basically you'd ask yourself, what number added to 1 will result in 0? Let's look in 4 bits, it will be easier. Remember we're limited only to 4 bits so any carry is dropped. $$ \begin{array}{lr} &0001\\ +&1111 \\ \hline &1 \boxed{0000} \end{array} $$ Likewise -2 will be the number that added to 2 will make zero so: $$ \begin{array}{lr} &0010\\ +&1110 \\ \hline &1 \boxed{0000} \end{array} $$so 1110 is -2, and so on. More generally you can say that to convert a number to its negative you have to first invert all the digits, because then we have: $$ \begin{array}{lr} &0101\\ +&1010 \\ \hline &1111 \end{array} $$And then by adding one to the inverted we'd get 0 instead of 1111 so: $$ \begin{array}{lr} &0101\\ +&1011 \\ \hline &1 \boxed{0000} \end{array} $$So the negative of 0101 (5 in decimal) is 1011 Notice that this is depended on the size of your register. so the negative of 000101 is 111011

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

You're welcome